happypotato:

Use the table to determine the location with the greatest amount of rainfall and the location with the least amount of rainfall. Using complete sentences, identify the location with the greatest amount of rainfall and the location with the least amount of rainfall, and express the ratio of their respective rainfalls as a whole number. Describe what this ratio tells you. (I can also send the table to you if you're willing to help)

happypotato:

1 attachment
jhonyy9:

bc. on the right side you see different length measures i think like a first step maybe more usefully to convert them in one the same measure what you think about this ?

jhonyy9:

do you know what is the SI. units measure for amount of rainfall ?

jhonyy9:

ok how i know is mm so you need convert every unit measures from the right side in mm do you know the way ?

happypotato:

no i'm not quite sure what youre talking about

jhonyy9:

for the first one 0,1173 m = ? mm

jhonyy9:

use this 1 m = 1000 mm

jhonyy9:

ok i will help you so using the rule of 3 simple we can write in this way 1 m mean 1000 mm 0,1173 m mean x mm and now cross multiplie and calcule the x value x = 0,1173×1000/1 = 0,1173×1000 = ? mm

jhonyy9:

how you multiplie the 0,1173 by 1000 ?

jhonyy9:

take the decimal 3 places on the right and will get 0,1173×1000 = 117,3 mm do you understand it now ?

jhonyy9:

ok for second one 11,4 m = ? mm

jhonyy9:

multiplie 11,4 by 1000 how many will get ?

jhonyy9:

11,4×1000 = ? take the decimal 3 places on the right so what will result ?

jhonyy9:

ok there wan be 11,400×1000 = ? so now is more easy

jhonyy9:

so look take the decimal 3 places on the right and will get 11,400×1000 = 11400 mm ok ?

jhonyy9:

the next one 25,6 cm = ? mm here you need to know that 1 cm = 10 mm

jhonyy9:

using this we can write 25,6×10=256 mm

jhonyy9:

the next two are given in mm so the last one please try convert it alone 480 cm = ? mm

jhonyy9:

@mayathegamergrl do you understand it ? do you can do it for medal ?

jhonyy9:

just you need multiplie the 480 by 10 480×10= ? mm

mayathegamergrl:

wait, thts it?

jhonyy9:

so = 4800 mm and now just you need to see what location has the greatest amount of rainfall

mayathegamergrl:

yh jus convert the cm to mm whcih you will multiply them by 10 each

jhonyy9:

1. has 117,3 mm 2. has 11400 mm 3. has 256 mm 4. has 57 mm 5. has 160 mm 6. has 4800 mm

mayathegamergrl:

yea

jhonyy9:

what has the greatest amount of rainfall ?

jhonyy9:

@happypotato do you understand the way ? questions ?

happypotato:

Yes I'm pretty sure i do. so it would be the second location right?

jhonyy9:

yeah

jhonyy9:

do you have questions ?

happypotato:

and what about the ratio ?

happypotato:

That was the second part of the question

happypotato:

Also thank you for being so helpful!!!

jhonyy9:

Using complete sentences, identify the location with the greatest amount of rainfall and the location with the least amount of rainfall, and express the ratio of their respective rainfalls as a whole number. Describe what this ratio tells you.

jhonyy9:

so you answered what is the greatest amount of rainfall what is the least amount ?

jhonyy9:

@jhonyy9 wrote:
1. has 117,3 mm 2. has 11400 mm 3. has 256 mm 4. has 57 mm 5. has 160 mm 6. has 4800 mm
look on this list

happypotato:

number 4

happypotato:

?

jhonyy9:

yes ofc

jhonyy9:

and express the ratio of their respective rainfalls as a whole number.

jhonyy9:

how many was the greatest ?

jhonyy9:

to find the ratio

happypotato:

11400 was the greatest

jhonyy9:

and the least ? to find the ratio you need divide the greatest by the least

happypotato:

57

happypotato:

and what does the ratio tell us?

jhonyy9:

so if how you see the location with the greatest amount of rainfall has 11400 mm and with the least has 57 mm what is you opinion what can this mean ? tell us ?

jhonyy9:

but you dont made the ratio calcule @happypotato

happypotato:

I'm not quite sure

jhonyy9:

@happypotato here is your question posted

happypotato:

thank you so much!

jhonyy9:

yw so now you understand it sure ?

jhonyy9:

you need calcule the ratio after this you get the result so you need use to write the answer on your problem posted

jhonyy9:

how much is the greatest : ?

happypotato:

number 2

jhonyy9:

11400 mm the least how much has ?

jhonyy9:

57 mm so now divide the greatest value with the least how many will get ?

jhonyy9:

11400 / 57 = ?

happypotato:

200

happypotato:

?

jhonyy9:

so you ve get the ratio is equal 200 what result from this ?

happypotato:

i dont understend your question?

jhonyy9:

the last question of your posted problem what is ?

jhonyy9:

Describe what this ratio tells you.

jhonyy9:

so from this ratio result that with this x number raining more in the greatest amount rainfall location then in the least amount rainfall location

jhonyy9:

this x is the ratio value

jhonyy9:

ok ? @extrinix can you please explain this little more understandably ? - tysm bc. i think @happypotato dont understand it clearly

Extrinix:

Yea, you have to remember happypotato is not a native english speaker.

jhonyy9:

@extrinix wrote:
Yea, you have to remember happypotato is not a native english speaker.
ok - tysm

jhonyy9:

so how i see this qc user go away is offline dont wait on your explain so sorry but tysm

jhonyy9:

choice 6 is 480 not 280 how many you get

jhonyy9:

and this 280 you say is in cm ,why ? first of all we need convert these results in one the same unit measure and how i remember the rainfall is measured in mm

Extrinix:

Needles, California Mt. Waialeale, Hawaii Puako, Hawaii Death Valley, California Mojave, California Hilo, Hawaii 0.1173m 11.4m 25.6cm 57mm 160mm 480cm Ok, so first off convert the chart I'm going to convert to centimeters because it's more accurate. 1mm = 0.1cm 1m = 100 cm 1 cm = 1 cm converted: Needles, California Mt. Waialeale, Hawaii Puako, Hawaii Death Valley, California Mojave, California Hilo, Hawaii 11.73cm 1140cm 25.6cm 5.7cm 16 cm 480cm now that that is converted, which location has the largest amount or rainfall and and which has the smallest amount of rainfall now to find the ratio, divide the largest number by the smallest number, \(\dfrac{1140}{5.7}\) so 200.

Extrinix:

It's not choice 6 specifically, its how many choices there are, and there are 6 total choices in the chart.

jhonyy9:

ohh ok

happypotato:

@extrinix wrote:
It's not choice 6 specifically, its how many choices there are, and there are 6 total choices in the chart.
So then what does the ratio tell me?

Extrinix:

@happypotato wrote:
@extrinix wrote:
It's not choice 6 specifically, its how many choices there are, and there are 6 total choices in the chart.
So then what does the ratio tell me?
the range of the numbers.

happypotato:

what do you mean by that?

happypotato:

@extrinix wrote:
@happypotato wrote:
@extrinix wrote:
It's not choice 6 specifically, its how many choices there are, and there are 6 total choices in the chart.
So then what does the ratio tell me?
the range of the numbers.
what do you mean by that?